OpenStudy (anonymous):

How can I make \(\sqrt{720}\) in simplified radical form?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

legendarysadist please help me more!

OpenStudy (anonymous):

Um, if you don't know then you can move along. XD You don't have to let me know that you don't know. :P

OpenStudy (anonymous):

Ok, there are a couple ways to go about this, but if you notice, \[\sqrt{36}*\sqrt{20}=\sqrt{720}\] From there you can also get \[\sqrt{4}*\sqrt{5}=\sqrt{20}\] So \[6*2*\sqrt{5}=\sqrt{720}\] 6*2=12 so the final result is\[12\sqrt{5}\]

OpenStudy (anonymous):

Wow! Thank you so much! I used to always be confused about how to get to simplified radical form, but you have made it seem very simple! Thank you for the clear explanation! Great help! :D

OpenStudy (anonymous):

@LegendarySadist ^

OpenStudy (anonymous):

Nope, 12 is the lowest you can go. We know this because we have simplified the square root side as far as it can go.